OpenStudy (anonymous):

WHAT IS 15% OF 23.60

OpenStudy (anonymous):

do you want 15% of the number cause thats 23.60 * .15 = 3.54

OpenStudy (anonymous):

WE JUST WENT OVER THIS STUFF! JUst let me find it in my book :D

OpenStudy (anonymous):

THANKS @ilikephysics2

OpenStudy (anonymous):

ok, nvm..

OpenStudy (anonymous):

that is simple (15/100)*23.60=0.15*23.60=3.54 is your answer

OpenStudy (anonymous):

I just showed it

OpenStudy (anonymous):

yeah, u just take the original price and multiply it by ur discount or % as a decimal and vwalalalala!!!